Shift report 20011009 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Oct 09 2001 - 15:43:56 EDT

  • Next message: Michael Murray: "Re: Allergies and breakdowns"

    ------------------------------------------------------------
    Date: Tue Oct  9 15:43:55 2001, Shift: 08:00-16:00
    Supervisor: Eun-Joo Kim, with: 
    Log book#9 pages 129 - 131
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam hours : about 2 & 1/2 hours
       "AGS Extraction power supply/magnet" message on TV
    
    2. Run 5264 : MRS@45, FS@12 
                  T3 HV  -> 20V lower than defalut value
                  T4-T5 HV -> default value
                  ~ 104 trigger 1
       Run 5265 : T3 HV  -> default value
                  T4-T5 HV -> 20V higher than defalut value  
                  ~ 91k trigger 1
       Run 5266 : same setting with Run 5265
                  FV set T3/T4/T5 threshold at 3.2V
                  ~ 224k trigger 1
    
       Pedestal(Camac & TPC), DVM Runs : 5267, 5268, 5269
    
    3. No new problems detected
     
    ---
    So far MRS@45(4B), FS@12(1/5A) : 711k trigger 1
     
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Oct 09 2001 - 15:44:42 EDT